# Hierarchical Learning and Computing over Space-Ground Integrated Networks

### v1

@@ I Introduction
With the rapid development of data sensing capabilities of smartphones, electric vehicles, and various sensors, a massive amount of generated data can be leveraged to train artificial intelligence (AI) models [1]. These AI models support various intelligent applications, such as autonomous driving [2] and environment monitoring [3]. Due to the advanced fifth-generation (5G) communication infrastructure in urban areas, Internet of Things (IoT) devices can quickly connect with cloud servers, providing necessary support for data transmission and AI model training. However, remote areas such as forests, deserts, and oceans, which have significant demand for intelligent services like disaster monitoring and deep-ocean exploration [4], lack coverage by ground base stations. This makes it challenging for data generated by IoT devices in these regions to be exploited for AI model training. To address this issue, an space-ground integrated network, serving as a complement and extension to terrestrial communication networks, leverages space-based communication and computing capabilities to provide wireless data transmission services for intelligent applications in remote areas [5, 6].
@@
The space-ground integrated networks comprise terrestrial IoT devices and the space computing power network (Space-CPN) [7]. Space-CPN typically includes unmanned aerial vehicles (UAVs), high-altitude platforms (HAPs), and various types of satellites. To accelerate the commercial deployment of space-ground communications, standards starting from 3GPP Release 17 specifications provide more support for non-terrestrial networks (NTN) [8]. Meanwhile, owing to the large-scale deployment of space-ground communication hardware in IoT devices, particularly the advancements in satellite communication hardware, direct connections between IoT devices and satellites have been successfully validated [9]. Traditional centralized learning requires transmitting data generated by widely distributed IoT devices to ground cloud servers for centralized processing via the space-ground integrated network, leading to substantial communication overhead and severe privacy concerns. To this end, a hierarchical learning and computing system, where remote IoT devices locally train AI models using their collected data, while the Space-CPN directly provides model aggregation and dissemination services for these IoT devices, can significantly enhance data transmission and learning efficiency [10, 11]. This paradigm can also be known as multi-tier federated learning (FL) [12] and FL with space-ground collaboration, as investigated in [13, 14, 15, 16].
@@
Satellites in Space-CPN can generally be classified into three categories based on their orbital altitudes, i.e., geostationary-earth-orbit (GEO) satellites, medium-earth-orbit (MEO) satellites, and low-earth-orbit (LEO) satellites [17]. To support the hierarchical learning and computing system, LEO mega-constellations, which are in active deployment nowadays (e.g., Starlink, Kuiper, Telesat, etc.) and are characterized by shorter communication delays and higher signal strengths, are highly suitable for effectively gathering local models from these remote IoT devices [18]. GEO satellites can efficiently manage and optimize communication paths between LEO satellites and are suitable for aggregating local models across the network and broadcasting the updated global model, due to their synchronous characteristics. Therefore, this paper considers a three-layer space-ground integrated network encompassing GEO satellites, LEO mega-constellations, and terrestrial edge devices for the implementation of learning and computing procedures, fully leveraging the computational capabilities of on-orbit processing and the advantages of inter-satellite communications [19, 20].
@@
In this three-layer space-ground integrated network, the energy consumption issue of the LEO satellites is particularly prominent. The power system of a satellite comprised of solar panels and battery cells, which generate electricity from sunlight and store the generated power, respectively. Due to the limitations imposed by the launch efficiency of rockets, LEO satellites must compromise in terms of size, weight, and power, with their capability of energy generation being particularly restricted [21]. Meanwhile, GEO satellites, during their orbital flight around the Earth, are predominantly exposed to sunlight, whereas LEO satellites frequently transition between daylight and nighttime conditions. This cyclical exposure poses significant challenges for power management in LEO satellites, since abuse of energy results in the rapid depletion of the satellite’s operational lifespan [22]. To aggregate the local models of ground IoT devices within LEO satellite networks through inter-satellite routing, in face of pronounced energy consumption concerns, it is imperative that the routing algorithms within LEO mega-constellations be designed to optimize for high energy efficiency and low power consumption [22, 23].
@@
Traditional routing problems in LEO satellite networks typically stem from providing relay connections for two distant terrestrial communication terminals. This can be modeled as an end-to-end shortest distance path problem, aiming to reduce the number of hops in the LEO satellite networks or to minimize the transmission delay or overhead [24, 25, 26]. Various methods can be applied to address these problems, including the Dijkstra algorithm, Bellman-Ford algorithm, and Floyd-Warshall algorithm [27]. These algorithms fundamentally model the satellite network as a graph and compute the shortest paths through iterative methods. However, in view of the hierarchical learning and computing architecture, LEO satellites that collect local models from ground devices are discretely distributed across the entire LEO mega-constellation. These LEO satellites can be regarded as terminals. On a large time scale, due to the Earth’s rotation and the mobility of LEO satellites, the identity of satellite terminals are different across communication rounds in the hierarchical learning procedures. On a smaller time scale, although intra-orbit communication topology can be considered stable [28], the inter-orbit communication topology changes rapidly [29]. This raises a critical challenge of efficiently aggregating these local models from terminals through inter-satellite routing with inter-satellite links (ISLs) in a highly dynamic network topology.
@@
In this paper, we consider a hierarchical learning and computing framework over a three-layer space-ground integrated network encompassing GEO satellites, LEO mega-constellations, and terrestrial edge devices for the implementation of learning and computing procedures. The aim of this paper is to design an energy-efficient global model aggregation routing algorithm for the hierarchical learning and computing framework based on the satellite network topology.
@@
    We propose a novel hierarchical learning and computing framework, tailed to space-ground integrated networks, to support collaboratively training on widely distributed terrestrial IoT devices with LEO mega-constellations and GEO satellites. In this space-ground integrated network, LEO satellites collect local models from terrestrial IoT devices and aggregate them through ISLs. GEO satellites is responsible for managing the routing paths and broadcasting the updated global model to these IoT devices.
@@
In realistic space-ground integrated network settings, extensive simulations on both Walker-Star and Walker-Delta constellations are conducted. The results demonstrate that our proposed TAEER scheme can achieve a better performance than other benchmarks, which verifies the superiority of the proposed algorithm.
@@ II System Model
In this section, we first present the network architecture of the space-ground integrated networks, as shown in Fig. 1. We then provide detailed information on the hierarchical learning and computing framework and the communication model.
@@ II-A1 Ground Networks
Consider a set $\mathcal{K}=\{1,\cdots,K\}$ of $K$ clusters located in different remote geographic regions, where the ground base stations and other terrestrial communication infrastructures are not available. In addition, a set $\mathcal{J}=\{1,\cdots,J\}$ of $J$ IoT devices are assumed to be distributed across these $K$ clusters, with each IoT device being located within a single cluster. We denote $\mathcal{K}_{j}$ as the set of IoT devices in cluster $j\in\mathcal{J}$, where $\cup_{j\in\mathcal{J}}\mathcal{K}_{j}=\mathcal{K}$. Each IoT device $i$ in cluster $j$, i.e., $i\in\mathcal{K}_{j}$, possesses its local dataset $\mathcal{D}_{i,j}=\{\bm{\theta}_{i,j}^{l},y_{i,j}^{l}\}_{i=1}^{|\mathcal{D}_{i,j}|}$. Let $\mathcal{D}_{j}=\cup_{i\in\mathcal{K}_{j}}\mathcal{D}_{i,j}$ denote the whole dataset in cluster $j$, and $\mathcal{D}=\cup_{j\in\mathcal{J}}\mathcal{D}_{j}$ as the overall dataset in the hierarchical learning system. By implementing hierarchical learning procedures, each IoT device participates in model training without sharing their local private data, utilizing their local computation capabilities.
@@ II-A2 Space Networks
To establish a hierarchical learning and computing framework covering all IoT devices in the remote clusters, we consider a LEO mega-constellation with a set of $\mathcal{N}=\{1,2,\dots,N\}$ orbit planes and three GEO satellites. The function of space networks is to provide model aggregation and model broadcasting services for ground IoT devices through inter-satellite routing. In orbit $n\in\mathcal{N}$, there are $S_{n}$ LEO satellites denoted as $\mathcal{S}_{n}=\left\{\mathsf{S}_{n,1},\mathsf{S}_{n,2},\dots,\mathsf{S}_{n,S_{n}}\right\}$ traveling around the Earth in the same direction. The total number of LEO satellites is thus given by $S=\sum_{n=1}^{N}S_{n}$. LEO satellites are responsible for servicing IoT devices within their respective clusters by facilitating hierarchical learning [30]. Additionally, these networks are responsible for the fusion of information across different clusters. Each terrestrial cluster can be supported by the LEO satellites from different orbits, or possibly the same orbit depending on their locations. However, due to the high mobility of LEO satellites, the LEO satellite serving a specific terrestrial cluster varies over time. Consequently, the concept of logical location is introduced to define the LEO satellite that serves a ground cluster at any point in time [25]. It is assumed that the entire Earth surface is covered by logic locations, which keep static and are taken over by the nearest satellite. This implies that once the coverage area of a given satellite no longer encompasses a particular terrestrial cluster, its service capabilities are handed over to the next satellite within the same orbit. By considering the service hand-off, the impact of satellite mobility on providing services for ground devices can be mitigated.
Figure 1: System model of a space-ground integrated network.
@@ II-B Hierarchical Learning and Computing Framework
In this subsection, we elaborate the hierarchical learning and computation framework over the space-ground integrated networks. First of all, the global objective is to minimize the global loss function defined as
$\min_{{\bm{x}}\in\mathbb{R}^{d}}f({\bm{x}}):=\sum_{j=1}^{J}\lambda_{j}f_{j}({\bm{x}}),$ (1)
where $\lambda_{j}\geq 0$ is the aggregation weight for IoT device $j$, $f_{j}(\cdot)$ is the local loss function of IoT device $j$, and ${\bm{x}}$ is the model parameter of dimension $d$. Based on the system model described in Section II-A, the global objective can be rewritten as
$f({\bm{x}})=\sum_{j=1}^{J}\sum_{i=1}^{|\mathcal{K}_{j}|}\lambda_{i,j}f_{i,j}\left({\bm{x}}\right),$ (2)
@@ II-B2 Global Aggregation
$\bm{x}^{t+1}:=\bm{x}^{t}+\sum_{j=1}^{J}\sum_{i=1}^{|\mathcal{K}_{j}|}\lambda_{i,j}\Delta\bm{x}_{i,j}^{t}.$ (6)
@@ II-C1 Ground-to-Space Link (GSL)
where ${}_{1}F_{1}(\cdot,\cdot,\cdot)$ denotes the hypergeometric function. The parameters $m$, $b_{0}$, and $\Omega$ signify the Nakagami fading coefficient, half of the scattered component’s average power, and the average power of the LOS component, respectively [32]. This expression can be closely approximated as a Gamma random variable by
@@
Besides the small-scale fading, the Free Space Loss due to the distance between the satellite and ground devices, can be modeled as $L=[\lambda/(4\pi l_{d})]^{2}$, with transmission distance $l_{d}$ and wavelength $\lambda$. Consequently, the channel coefficient of the communication link between IoT devices and the LEO satellite is $|H|^{2}=L|h|^{2}$.
@@ III-A Characterization of Space Topology
This indicates that the distance between two satellites in different orbits should not exceed the minimum communication range of the satellites if they want to establish connection with each other. In a nutshell, the GEO satellites maintain awareness of the edges that can be established by ISLs in each snapshot. This forms a snapshot table, which can be accessed and reviewed by the GEO satellites at any time.

### v2

@@ I Introduction
With the rapid development of data sensing capabilities of smartphones, electric vehicles, and various sensors, a massive amount of generated data can be leveraged to train artificial intelligence (AI) models [1]. These AI models support various intelligent applications, such as autonomous driving [2] and environment monitoring [3]. Due to the advanced fifth-generation (5G) communication infrastructure in urban areas, Internet of Things (IoT) devices can quickly connect with cloud servers, providing necessary support for data transmission and AI model training. However, remote areas such as forests, deserts, and oceans, which have significant demand for intelligent services like disaster monitoring and deep-ocean exploration [4], lack coverage by ground base stations. This makes it challenging for data generated by IoT devices in these regions to be exploited for AI model training. To address this issue, an space-ground integrated network (SGIN), serving as a complement and extension to terrestrial communication networks, leverages space-based communication and computing capabilities to provide wireless data transmission services for intelligent applications in remote areas [5, 6]. The SGINs comprise terrestrial IoT devices and the space computing power network (Space-CPN) [7]. Space-CPN typically includes unmanned aerial vehicles (UAVs), high-altitude platforms (HAPs), and various types of satellites.
@@
To accelerate the commercial deployment of space-ground communications, standards starting from 3GPP Release 17 specifications provide more support for non-terrestrial networks (NTN) [8]. Meanwhile, owing to the large-scale deployment of space-ground communication hardware in IoT devices, particularly the advancements in satellite communication hardware, direct connections between IoT devices and satellites have been successfully validated [9]. Traditional centralized learning requires transmitting data generated by widely distributed IoT devices to ground cloud servers for centralized processing via the SGIN, leading to substantial communication overhead and severe privacy concerns. To this end, a hierarchical learning and computing system, where remote IoT devices locally train AI models using their collected data, while the Space-CPN directly provides model aggregation and dissemination services for these IoT devices, can significantly enhance data transmission and learning efficiency [10, 11]. This paradigm can also be known as multi-tier federated learning (FL) [12]. In the context of existing research on learning and computing in SGINs, FL in SGINs can be generally classified into two distinct categories. The first involves terrestrial devices performing local model training while satellite-based systems handle model aggregation (e.g., [13, 14, 15]), whereas the second employs space-based nodes for local model training with subsequent aggregation conducted at ground stations (e.g., [16, 17]). Moreover, recent literature also studied edge computing techniques in SGINs, particularly focusing on critical aspects including service caching mechanisms [18], task offloading strategies [19], and on-orbit resource allocation schemes [20].
@@
Satellites in Space-CPN can generally be classified into three categories based on their orbital altitudes, i.e., geostationary-earth-orbit (GEO) satellites, medium-earth-orbit (MEO) satellites, and low-earth-orbit (LEO) satellites [21]. To support the hierarchical learning and computing system, LEO mega-constellations, which are in active deployment nowadays (e.g., Starlink, Kuiper, Telesat, etc.) and are characterized by shorter communication delays and higher signal strengths, are highly suitable for effectively gathering local models from these remote IoT devices [22]. GEO satellites can efficiently manage and optimize communication paths between LEO satellites and are suitable for aggregating local models across the network and broadcasting the updated global model, due to their synchronous characteristics. Therefore, this paper considers a three-layer SGIN encompassing GEO satellites, LEO mega-constellations, and terrestrial edge devices for the implementation of learning and computing procedures, fully leveraging the computational capabilities of on-orbit processing and the advantages of inter-satellite communications [23, 24]. In terrestrial networks, the scheduling of communication schemes for distributed model training, such as data, model, and pipeline parallelism in AI center networks, is critically important [25]. While these methods may not significantly affect the final performance of model training, they can greatly accelerate the overall training process. For instance, communication patterns between servers and nodes, such as gather, scatter, and all-reduce, are specifically designed for different types of computational tasks. Therefore, designing a communication scheme tailored to AI traffic for a three-layer SGIN, specifically a routing method for model aggregation in FL procedures, is highly meaningful.
@@
In this three-layer SGIN, the energy consumption issue of the LEO satellites is particularly prominent. The power system of a satellite comprised of solar panels and battery cells, which generate electricity from sunlight and store the generated power, respectively. Due to the limitations imposed by the launch efficiency of rockets, LEO satellites must compromise in terms of size, weight, and power, with their capability of energy generation being particularly restricted [26]. Meanwhile, GEO satellites, during their orbital flight around the Earth, are predominantly exposed to sunlight, whereas LEO satellites frequently transition between daylight and nighttime conditions. This cyclical exposure poses significant challenges for power management in LEO satellites, since abuse of energy results in the rapid depletion of the satellite’s operational lifespan [27]. Specifically, the batteries of the satellites have inherent limitations in terms of recharge/discharge cycles, commonly known as the depth of discharge (DOD) cycle. These limitations restrict not only the useful life of the batteries but also that of the satellites [28, 29, 30]. Therefore, to aggregate the local models of ground IoT devices within LEO satellite networks through inter-satellite routing, in face of pronounced energy consumption concerns, it is imperative that the routing algorithms within LEO mega-constellations be designed to optimize for high energy efficiency and low power consumption [27, 31].
@@
Compared to conventional terrestrial routing algorithms, routing in LEO satellite networks faces three major challenges: highly dynamic topologies, constrained on-board resources, and complex inter-satellite link (ISL) conditions. Existing studies have primarily addressed the above challenges in scenarios focused on providing relay connections for two distant terrestrial communication terminals [32]. This can be modeled as an end-to-end shortest distance path problem, aiming to reduce the number of hops in the LEO satellite networks or to minimize the transmission delay or overhead [33, 34, 35]. Various methods can be applied to address these problems, including the Dijkstra algorithm, Bellman-Ford algorithm, and Floyd-Warshall algorithm [36]. These algorithms fundamentally model the satellite network as a graph and compute the shortest paths through iterative methods. However, in view of the hierarchical learning and computing architecture, LEO satellites that collect local models from ground devices are discretely distributed across the entire LEO mega-constellation. These LEO satellites can be regarded as terminals. On a large time scale, due to the Earth’s rotation and the mobility of LEO satellites, the identity of satellite terminals are different across communication rounds in the hierarchical learning procedures. On a smaller time scale, although intra-orbit communication topology can be considered stable [37], the inter-orbit communication topology changes rapidly [38]. This raises a critical challenge of efficiently aggregating these local models from terminals through inter-satellite routing with ISLs in a highly dynamic network topology.
@@
In this paper, we consider a hierarchical learning and computing framework over a three-layer SGIN encompassing GEO satellites, LEO mega-constellations, and terrestrial edge devices for the implementation of learning and computing procedures. The aim of this paper is to design an energy-efficient global model aggregation routing algorithm for the hierarchical learning and computing framework based on the satellite network topology.
@@
    We propose a novel hierarchical learning and computing framework, tailed to SGINs, to support collaboratively training on widely distributed terrestrial IoT devices with LEO mega-constellations and GEO satellites. In this SGIN, LEO satellites collect local models from terrestrial IoT devices and aggregate them through ISLs. GEO satellites is responsible for managing the routing paths and broadcasting the updated global model to these IoT devices.
@@
In realistic SGIN settings, extensive simulations on both Walker-Star and Walker-Delta constellations are conducted. The results demonstrate that our proposed TAEER scheme can achieve a better performance than other benchmarks, which verifies the superiority of the proposed algorithm.
@@ II System Model
In this section, we first present the network architecture of the SGINs, as shown in Fig. 1. We then provide detailed information on the hierarchical learning and computing framework and the communication model.
@@ II-A1 Ground Networks
Consider a set $\mathcal{K}=\{1,\cdots,K\}$ of $K$ clusters located in different remote geographic regions, where the ground base stations and other terrestrial communication infrastructures are not available. In addition, a set $\mathcal{J}=\{1,\cdots,J\}$ of $J$ IoT devices are assumed to be distributed across these $K$ clusters, with each IoT device being located within a single cluster. We denote $\mathcal{K}_{j}$ as the set of IoT devices in cluster $j\in\mathcal{K}$, where $\cup_{j\in\mathcal{K}}\mathcal{K}_{j}=\mathcal{K}$. Each IoT device $i$ in cluster $j$, i.e., $i\in\mathcal{K}_{j}$, possesses its local dataset $\mathcal{D}_{i,j}=\{\bm{\theta}_{i,j}^{l},y_{i,j}^{l}\}_{i=1}^{|\mathcal{D}_{i,j}|}$. Let $\mathcal{D}_{j}=\cup_{i\in\mathcal{K}_{j}}\mathcal{D}_{i,j}$ denote the whole dataset in cluster $j$, and $\mathcal{D}=\cup_{j\in\mathcal{J}}\mathcal{D}_{j}$ as the overall dataset in the hierarchical learning system. By implementing hierarchical learning procedures, each IoT device participates in model training without sharing their local private data, utilizing their local computation capabilities. Specific application scenarios include large-scale geospatial applications such as environmental monitoring, ocean exploration, and smart agriculture.
@@ II-A2 Space Networks
To establish a hierarchical learning and computing framework covering all IoT devices in the remote clusters, we consider a LEO mega-constellation with a set of $\mathcal{N}=\{1,2,\dots,N\}$ orbit planes and three GEO satellites. The function of space networks is to provide model aggregation and model broadcasting services for ground IoT devices through inter-satellite routing. In orbit $n\in\mathcal{N}$, there are $S_{n}$ LEO satellites denoted as $\mathcal{S}_{n}=\left\{\mathsf{S}_{n,1},\mathsf{S}_{n,2},\dots,\mathsf{S}_{n,S_{n}}\right\}$ traveling around the Earth in the same direction. The total number of LEO satellites is thus given by $S=\sum_{n=1}^{N}S_{n}$. To leverage the low latency and wide coverage characteristics of LEO satellite networks to provide access services for IoT devices lacking ground communication infrastructure, LEO satellites are tasked with servicing IoT devices within their respective clusters by enabling hierarchical learning [Le2024survey]. Additionally, these networks are responsible for the fusion of information across different clusters. Each terrestrial cluster can be supported by the LEO satellites from different orbits, or possibly the same orbit depending on their locations. However, due to the high mobility of LEO satellites, the LEO satellite serving a specific terrestrial cluster varies over time. Consequently, the concept of logical location is introduced to define the LEO satellite that serves a ground cluster at any point in time [34]. It is assumed that the entire Earth surface is covered by logic locations, which keep static and are taken over by the nearest satellite. This implies that once the coverage area of a given satellite no longer encompasses a particular terrestrial cluster, its service capabilities are handed over to the next satellite within the same orbit. By considering the service hand-off, the impact of satellite mobility on providing services for ground devices can be mitigated.
Figure 1: System model of a SGIN.
@@ II-B Hierarchical Learning and Computing Framework
In this subsection, we elaborate the hierarchical learning and computation framework over the SGINs. First of all, the global objective is to minimize the global loss function defined as
$\min_{{\bm{x}}\in\mathbb{R}^{d}}f({\bm{x}}):=\sum_{n=1}^{J}\lambda_{n}f_{n}({\bm{x}}),$ (1)
where $\lambda_{n}\geq 0$ is the aggregation weight for IoT device $n$, $f_{n}(\cdot)$ is the local loss function of IoT device $n$, and ${\bm{x}}$ is the model parameter of dimension $d$. Based on the system model described in Section II-A, the global objective can be rewritten as
$f({\bm{x}})=\sum_{j=1}^{K}\sum_{i=1}^{|\mathcal{K}_{j}|}\lambda_{i,j}f_{i,j}\left({\bm{x}}\right),$ (2)
@@ II-B2 Global Aggregation
$\bm{x}^{t+1}:=\bm{x}^{t}+\sum_{j=1}^{K}\sum_{i=1}^{|\mathcal{K}_{j}|}\lambda_{i,j}\Delta\bm{x}_{i,j}^{t}.$ (6)
Global aggregation is a typical procedure in model training that integrates communication and computation. Designing an efficient model aggregation routing method for the hierarchical learning and computing system proposed in this paper is the key to improving the efficiency of model training.
@@ II-C1 Ground-to-Space Link (GSL)
where ${}_{1}F_{1}(\cdot,\cdot,\cdot)$ denotes the hypergeometric function. The parameters $m$, $b_{0}$, and $\Omega$ signify the Nakagami fading coefficient, half of the scattered component’s average power, and the average power of the LOS component, respectively [40]. The physical meanings of the hypergeometric function lies in the fact that when the LOS component is dominant, the contribution of the hypergeometric function becomes significant, capturing the enhancement effect of the LOS signal. Conversely, when the LOS component is weak, the impact of the hypergeometric function diminishes, reflecting the signal characteristics dominated by multipath scattering. This expression can be closely approximated as a Gamma random variable by
@@
Besides the small-scale fading, the Free Space Loss due to the distance between the satellite and ground devices, can be modeled as $L=[\lambda/(4\pi l_{d})]^{2}$, with transmission distance $l_{d}$ and wavelength $\lambda$. Consequently, the channel coefficient of the communication link between IoT devices and the LEO satellite is $|H|^{2}=L|h|^{2}$. We assume an uplink method based on frequency-division multiple access (FDMA) to avoid intra-cell interference and ignore the inter-cell interference signals at the receiver for simplification.
@@ III-A Characterization of Space Topology
Figure 3: Distance constraint for inter-orbit laser ISLs.
@@
This indicates that the distance between two satellites in different orbits should not exceed the minimum communication range of the satellites if they want to establish connection with each other, as illustrated in Fig. 3. The network topology constraints are defined as follows. While distance constraints may be satisfied, two satellites within the same orbit cannot establish a link if there is an intermediate satellite between them, i.e., $(\mathsf{S}_{n,i},\mathsf{S}_{n,j})\in\mathcal{E}(m),|i-j|=1,|i-j|=S_{n}-1$. For satellites in different orbits, when multiple satellites in another orbit satisfy the distance constraint, only the link with the nearest satellite is considered. In a nutshell, the GEO satellites maintain awareness of the edges that can be established by ISLs in each snapshot. This forms a snapshot table, which can be accessed and reviewed by the GEO satellites at any time.